# Federated Hierarchical Reinforcement Learning for Adaptive Traffic Signal Control

### v1

@@ 5.3 Real-world Experiments
In addition to the synthetic examples, we also apply the proposed algorithms to real-world networks near Columbia University, as shown in Fig. 15. We configure the traffic demand using the methodology in Section  5.1 to calibrate the simulation. The results are analyzed in terms of travel time, waiting time, and clustering patterns.
@@
5.3.1 Reward Convergence
Fig. 16 presents the convergence curves of the reward. The results indicate that all other algorithms achieve convergence. The decentralized method has the lowest reward. Although the centralized method achieves the highest training reward, our proposed methods outperform it in evaluation tests, as discussed in the next section.
@@ 5.3.2 Performance Comparison
Fig. 17 and 18 show the average travel and waiting time in a real-world traffic network, where FedFomoLight achieves the best performance with both metrics. FedClusterLight and FedAvg perform similarly, slightly worse than FedFomoLight, but still significantly better than the other methods. A more detailed comparison is evident in the zoomed-in area. In contrast, the decentralized method performs the worst, exhibiting the highest travel time, while the centralized method also performs poorly. The results indicate that although the proposed methods may not surpass the centralized framework in synthetic networks, they exhibit significantly better performance in real-world networks. This is likely due to the complexity of real-world networks, where the centralized method may not effectively capture the diverse characteristics of intersections within the hierarchical structure.
Figure 17: Average Travel Time for Real World Traffic Flow
Figure 18: Average Waiting Time for Real World Traffic Flow
Table. 6 presents the estimated communication cost, calculated using the method in Sec. 5.1.2. The results presented in the table demonstrate that our proposed algorithms, FedClusterLight and FedFomoLight, achieve relatively lower communication costs than the centralized method and FedAvg.
@@ 5.3.3 Sensitivity Analysis
We analyzed the clustering results from synthetic experiments and applied a similar approach to real-world experiments. Figure 19 shows the cluster distribution produced by FedClusterLight after training. The results indicate that FedClusterLight can capture some adjacency relationships between certain intersections; however, the overall pattern in the relaying process remains weak. Fig. 20 presents the hierarchical clustering results of FedFomoLight after training, obtained from a parameter-compression-based matrix. FedFomoLight identifies correlations between model parameters and the traffic network topology, grouping intersection agents accordingly. Cluster group 1 (Fig. 20) includes intersections along Amsterdam Avenue and Manhattan Avenue. Cluster Group 2 (Fig. 20) consists of intersections along Broadway, while Cluster Group 3 (Fig. 20) represents the intersection at the network’s corner. Additionally, we randomly selected three intersections and visualized their six most similar counterparts. The results indicate that FedFomoLight captures adjacency relationships among intersections and highlight similarities in their structural features.
Figure 19: Cluster pattern of FedClusterLight in real-world network
@@ 6 CONCLUSION
These results demostrate the practical benefits of HFRLs: reduced communication overhead, improved scalability for large-scale urban systems, and greater flexibility in handling diverse traffic conditions. Future studies could incorporate the following directions:
@@
Acknowledgements

### v2

@@ 5.3 Real-world Experiments
In addition to the synthetic examples, we also apply the proposed algorithms to real-world networks near Columbia University, as shown in Fig. 12. We configure the traffic demand using the methodology in Section 5.1 to calibrate the simulation. The results are analyzed in terms of travel time, waiting time, and clustering patterns.
@@
5.3.1 Reward Stabilization
Fig. 13 presents the reward stabilization curves. The training rewards of all algorithms reach a stable plateau within the fixed number of episodes, with only minor oscillations near the stabilized stage. Such small fluctuations are common in FRL due to heterogeneous local environments and asynchronous updates (Qi2021; Xie2022; Emara2024). The decentralized method has the lowest reward. Our proposed methods outperform other methods in evaluation tests, as discussed in the next section.
@@ 5.3.2 Performance Comparison
Tab. 6 shows the average travel and waiting time in a real-world traffic network, where FedFomoLight achieves the best performance with both metrics. FedClusterLight and FedAvg perform similarly, slightly worse than FedFomoLight, but still significantly better than the other methods. In contrast, the decentralized method performs the worst, exhibiting the highest travel time, while the centralized method also performs poorly. The results indicate that although the proposed methods may not surpass the centralized framework in synthetic networks, they exhibit significantly better performance in real-world networks. This is likely due to the complexity of real-world networks, where the centralized method may not effectively capture the diverse characteristics of intersections within the hierarchical structure. Tab. 11 presents the estimated communication cost, calculated using the method in Sec. 5.1.2. The results presented in the Tab. 11 demonstrate that our proposed algorithms, FedClusterLight and FedFomoLight, achieve relatively lower communication costs than the centralized method and FedAvg. Although the decentralized method eliminates the need for communication with a central server, it performs worse than the other methods in real-world settings, as shown in Table 6. As a result, more vehicles remain active on the road, leading to non-negligible communication overhead. This outcome differs from the synthetic experiments, in which the decentralized method exhibits competitive performance and correspondingly lower communication costs.
@@ 5.3.3 Sensitivity Analysis
We conducted a sensitivity analysis on the real-world network and clarified how to select the best $K$ for different networks. Fig. 14 shows the cluster distributions produced by FedClusterLight after training with different numbers of clusters. The results indicate that FedClusterLight can capture adjacency relationships between certain intersections. The main groups remain quite stable across different $K$ values, such as intersections at the corners or those on the main road. As $K$ increases, some new groups may emerge, which aligns with our intuition. To select the optimal $K$ for training, we conducted a preliminary experiment, as shown in Fig. 15, which indicates that the algorithm achieves its best performance when $K=4$.
Figure 14: Cluster Patterns of FedClusterLight in the Real-World Network with Different $K$
Figure 15: Reward Curve of FedClusterLight with Different Numbers of Clusters $K$
Fig. 16 presents the hierarchical clustering results of FedFomoLight after training, obtained from a parameter-compression-based matrix. FedFomoLight identifies correlations between model parameters and the traffic network topology, grouping intersection agents accordingly. Cluster group 1 (Fig. 16) includes intersections along Amsterdam Avenue and Manhattan Avenue. Cluster Group 2 (Fig. 16) consists of intersections along Broadway, while Cluster Group 3 (Fig. 16) represents the intersection at the network’s corner. Additionally, we randomly selected three intersections and visualized their six most similar counterparts. The results indicate that FedFomoLight captures adjacency relationships among intersections and highlight similarities in their structural features.
@@ 6 CONCLUSION
These results demonstrate the practical benefits of HFRLs: reduced communication overhead, improved scalability for large-scale urban systems, and greater flexibility in handling diverse traffic conditions. Future studies could incorporate the following directions:
@@
  4. (4)
    Incorporate fairness-aware learning objectives into the HFRL framework. Future studies could design reward functions or aggregation strategies that ensure intersections receive equally effective models, preventing systematic bias where certain locations experience persistently longer delays.
Acknowledgments